A new municipal refuse-collection truck can be purchased for $84,000. Its expected useful life is six years, at which time its market value will be zero. Annual receipts less expenses will be approximately $18,000 per year over the six-year study period. Use the PW method and a MARR of 18% to determine whether this is a good investment.


1
Expert's answer
2021-10-05T08:02:02-0400

PW

= - C0 + PW of annual receipt less expenses as annuity

"= - C0 + B * PVAF (n, r)\\

= - 84,000 + 18,000 (6, 18 \%)\\

= - 84,000 + 18,000 / 18 \% \times [1 - (1 + 18 \%)-6]\\

= -21,043.15\\"

Since the PW is negative, this is not a good investment.